OpenStudy (anonymous):

Solve this absolute value equation and show all work. 9 + |2x - 3| = 7

OpenStudy (anonymous):

isolate the absolute value

OpenStudy (anonymous):

You want to get the absolute value on one side by itself. |2x-3|=7-9 Do you see a problem here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Tell me whats wrong

OpenStudy (anonymous):

Can a absolute value be negative?

OpenStudy (anonymous):

|2x-3| = -2 Is this right?

OpenStudy (anonymous):

absolute values are always positive. So whats your solution set?

OpenStudy (anonymous):

S = the empty set

OpenStudy (anonymous):

There is no solution